# On the cosmological abundance of magnetic monopoles

### v1

@@ 1 Introduction
Magnetic monopole is one of the most fascinating objects in theoretical physics Preskill:1984gd ; Stone:1984vom ; Groom:1986ps ; Giacomelli:2000de ; Shnir:2005vvi ; Balestra:2011ks ; Rajantie:2012xh ; Patrizii:2015uea ; Patrizii:2019eud ; Mavromatos:2020gwk . It stems from the natural requirement of extending the symmetry between electric and magnetic fields beyond the source-free Maxwell’s equations. Its consistency with quantum mechanics is demonstrated in 1931 by Dirac, which has led to an intriguing explanation for the charge quantization observed in nature Dirac:1931kp . In 1974 ’t Hooft and Polyakov demonstrated that magnetic monopoles necessarily appear as topological defects in particle theories of grand unification tHooft:1974kcl ; Polyakov:1974ek , with the topological charge identified with the magnetic charge. It was soon realized that the cosmological abundance of magnetic monopoles poses a serious problem for such scenarios Preskill:1979zi ; Vilenkin:2000jqa ; Weinberg:2012pjx , the solution of which gave a strong boost to the birth of the inflation theory Starobinsky:1980te ; Guth:1980zm . Because magnetic monopole is deeply connected with the fundamental symmetries of nature at very high energy, the discovery of even a single magnetic monopole would have far-reaching implications. It is thus intriguing that the recent Amaterasu cosmic ray particle found by the Telescope Array Group TelescopeArray:2023sbd might be interpreted as a magnetic monopole with extremely high energy Cho:2023krz ; Frampton:2024shp .
@@
Although it is possible that cosmic inflation has diluted away magnetic monopoles so that they end with an abundance too small to have observable consequences, it is not necessarily the case Preskill:1984gd ; ParticleDataGroup:2022pth . Relatively light magnetic monopoles might be produced after inflation through various mechanisms Kibble:1976sj ; Zurek:1985qw ; Murayama:2009nj ; Das:2021wei ; Kobayashi:2021des ; Kobayashi:2022qpl ; He:2022wwy so that their abundance is not affected by the dilution. Even if for superheavy magnetic monopoles that are indeed affected by inflation, their final abundance still depends on its initial value and the actual amount of dilution and which can be quite model-dependent and not well-constrained Epele:2007ic ; Senoguz:2015lba ; Kephart:2017esj ; Lazarides:2019xai ; Chakrabortty:2020otp ; Vento:2020vsq . Alternative scenarios to inflation could also be envisioned Brandenberger:2009jq . Therefore we choose to be open-minded and inquire what model-independent constraints can be obtained on the cosmological monopole abundance, which is taken as a free parameter determined by some unknown production mechanism in the very early universe. Here by “model-independent constraints” we refer to constraints that can be derived from the electromagnetic and gravitational properties of the magnetic monopole only. Other properties that depend on the model detail (such as the catalysis of baryon number violation Rubakov:1982fp ; Callan:1982ah ; Callan:1982au ) will not be assumed.
@@
In this work we shall however demonstrate that the Debye shielding bound suggested by ref. Medvedev:2017jdn cannot be used to obtain meaningful bounds on the cosmological monopole abundance. The reason is that, as we shall see, the concept of Debye shielding simply does not hold for magnetic monopole plasma in the presence of magnetic fields that are not curl-free. A similar issue in the context of solar physics was discussed in Appendix B of ref. Spicer:1984uui , where the authors commented that “Debye shielding will not shield out a transverse electric field.”
For non-relativistic magnetic monopoles that do not catalyze baryon number violation, the existing bounds on their cosmological abundance then become quite limited. We revisit the mass density constraint, the bound from direct search in cosmic rays and the conventional Parker bounds. We found that the Andromeda Parker bound which was once proposed to constrain the abundance of magnetic black holes Bai:2020spd , can be employed to constrain the abundance of ordinary magnetic monopoles effectively.
@@ 2.3 Anatomy of the magnetic monopole plasma with a two-fluid model
In order to present a solid argument, let us show that the assumption that Debye shielding is achieved in a magnetic monopole plasma with test electric currents leads to inconsistencies, and thus cannot be used to obtain a meaningful constraint on the cosmological monopole abundance. We start by considering the magnetic monopole plasma using a two-fluid model, with one component fluid composed of magnetic monopoles, and the other component fluid composed of magnetic anti-monopoles. We may then write down the Euler equations for the magnetic monopole fluid and the anti-monopole fluid:
@@
in which v is the fluid velocity associated with the electric currents333v should be thus related to $\textbf{J}_{e}$ via an equation like Eq. (18), with an additional equation for the continuity of the electric charge. These details are not needed in the following so we do not bother to present them explicitly., and $\sigma$ denotes the electric conductivity which may be a function of the position and time. Finally, there should be equations of state which express the pressure as a function of the number density and temperature:
@@
Similar assumptions are made in the usual treatment of electron-proton plasma in a two-fluid model Thorne:2017mcp . In particular, the first assumption implies that we are effectively working at scales much larger than the mean separation between magnetic monopoles and the magnetic monopole plasma is not very dense so that the monopole annihilation can be neglected.
@@ 3 Revisiting current bounds on the cosmological monopole abundance
    Parker bound.
All these bounds have been discussed to various extent in previous literature Burdin:2014xma ; Patrizii:2015uea ; Patrizii:2019eud ; Mavromatos:2020gwk ; ParticleDataGroup:2022pth ; Kobayashi:2023ryr . The emphasis of the discussion here is to examine whether magnetic monopoles may account for all (or the majority) of the dark matter, and the assumptions and uncertainties involved in the analysis.
Bounds on the cosmological monopole abundance can be represented in various manners. In many cases one considers magnetic monopoles (including anti-monopoles) populated with the rest number density $n_{M}$ and an isotropic average velocity $v$ in some volume defined in a reference frame $S$. Here the reference frame and the volume depend on the type of local measurement under consideration. For example, for direct search in cosmic rays, the reference frame is fixed on the Earth and the volume is associated with the size of the solar system, while for the Parker bound from the galactic magnetic field, the reference frame is fixed on the Milky way galaxy and the volume is associated with the galactic magnetic field region in the Milky way. The local isotropic flux of magnetic monopoles is defined as
$F_{M}=\frac{\gamma n_{M}v}{4\pi},$ (43)
where we have introduced $\gamma\equiv(1-v/c)^{-1/2}$ which is the Lorentz boost factor of the magnetic monopoles measured in the reference frame $S$ associated with the local measurement under consideration. Eq. (43) means that if one considers an area element $dS$ with its normal direction n, then during a short time $dt$ the number of monopoles (and anti-monopoles) that cross $dS$ with velocity direction in a small solid angle $d\Omega$ around n is given by $F_{M}dtdSd\Omega$. Therefore the unit of $F_{M}$ is written as $\text{cm}^{-2}\text{s}^{-1}\text{sr}^{-1}$. To gain an intuitive feeling of the flux, we may compute the mean separation $d_{M}=n_{M}^{-1/3}$ of magnetic monopoles as a function of $F_{M}$ and $v$ as
$d_{M}\simeq 288\,\text{km}\times\left(\frac{\gamma v}{10^{-3}c}\right)^{1/3}\left(\frac{F_{M}}{10^{-16}\text{cm}^{-2}\text{s}^{-1}\text{sr}^{-1}}\right)^{-1/3},$ (44)
@@
In order to discuss the conversion between $F_{M}$ and $\Omega_{M}$, we introduce $\rho_{M}=n_{M}mc^{2}\gamma$ which is the energy density of the magnetic monopoles in the $S$ frame. The conversion between $F_{M}$ and $\Omega_{M}$ is then given by
$F_{M}$ $=\frac{\rho_{\text{crit}}\Delta v}{4\pi mc^{2}}\Omega_{M},$ (47)
$\Omega_{M}$ $=\frac{4\pi mc^{2}}{\rho_{\text{crit}}\Delta v}F_{M},$ (48)
@@
The conversion between $F_{M}$ and $\Omega_{M}$ critically depends on the velocity $v$ of the magnetic monopoles. For monopoles clustered with the galaxy ($m\gtrsim 10^{18}\,\text{GeV}/c^{2}$), $v\simeq 10^{-3}c$ which is the virial velocity of the galaxy. Due to the presence of galactic and intracluster magnetic fields lighter magnetic monopoles can be accelerated to higher (and even relativistic) velocities. It is estimated that magnetic monopoles with masses around $10^{16}\sim 10^{17}\,\text{GeV}$ can be accelerated to $v\simeq 10^{-2}c$ Medvedev:2017jdn ; ParticleDataGroup:2022pth which is the escape velocity of large galaxy clusters Medvedev:2017jdn , while magnetic monopoles with masses smaller than $10^{11}\sim 10^{13}\,\text{GeV}$ are expected to be in the relativistic regime Wick:2000yc ; Burdin:2014xma ; Medvedev:2017jdn ; ParticleDataGroup:2022pth . The dependence of the average monopole velocity on its mass is associated with its complicated acceleration history Wick:2000yc ; Perri:2023ncd and thus cannot be predicted precisely.
To facilitate comparison between different constraints, let us express Eqs. (47) and (48) using various benchmark values
@@
Here for $B$ we consider only the regular magnetic field Fletcher:2003ec . The Parker bound from magnetic fields of the Andromeda galaxy was employed by ref. Bai:2020spd ; Ghosh:2020tdu to constrain the abundance of extremal magnetic black holes. It was then pointed out in ref. Kobayashi:2023ryr that whether such magnetic black holes remain clustered with the Andromeda galaxy is uncertain. Here we use the Andromeda version of the Parker bound to constrain the abundance of ordinary magnetic monopoles with $g=g_{D}$. The boundary between clustered and unclustered monopoles in the case of the Andromeda galaxy is expected to be at $m\simeq 10^{19}\,\text{GeV}/c^{2}$ Kobayashi:2023ryr , again with a transition range spanning $2\sim 3$ orders of magnitude. Therefore somewhat large uncertainty on the overdensity $\Delta$ is involved if one wants to convert the Andromeda Parker bound on the monopole flux at large monopole mass $m\sim 10^{19}\,\text{GeV}/c^{2}$ to a corresponding bound on the monopole density parameter. For much lighter magnetic monopoles with $m\lesssim 10^{19}\,\text{GeV}/c^{2}$ it is likely that they are unclustered with $\Delta\simeq 1$.
@@
$l_{c}\lesssim\frac{v}{2g}\left(\frac{m\gamma v}{F_{M}}\right)^{1/2}.$ (65)
Using benchmark values, Eq. (65) can be expressed as [^5] To obtain the numbers, we use $g_{D}\simeq 8.2\times 10^{-7}\,\text{GeV}^{1/2}\,\text{cm}^{1/2}$ in Gaussian units, which can be derived from the Dirac quantization condition $eg_{D}=\hbar c/2$ and $\alpha=\frac{e^{2}}{\hbar c}$. (we approximate $\gamma\simeq 1$)

### v3

@@ 1 Introduction
Magnetic monopole is one of the most fascinating objects in theoretical physics Preskill:1984gd ; Stone:1984vom ; Groom:1986ps ; Giacomelli:2000de ; Shnir:2005vvi ; Balestra:2011ks ; Rajantie:2012xh ; Patrizii:2015uea ; Patrizii:2019eud ; Mavromatos:2020gwk . It stems from the natural requirement of extending the symmetry between electric and magnetic fields beyond the source-free Maxwell’s equations. Its consistency with quantum mechanics is demonstrated in 1931 by Dirac, which has led to an intriguing explanation for the charge quantization observed in nature Dirac:1931kp . In 1974 ’t Hooft and Polyakov demonstrated that magnetic monopoles necessarily appear as topological defects in particle theories of grand unification tHooft:1974kcl ; Polyakov:1974ek , with the topological charge identified with the magnetic charge. It was soon realized that the cosmological abundance of magnetic monopoles poses a serious problem for such scenarios Zeldovich:1978wj ; Preskill:1979zi ; Vilenkin:2000jqa ; Weinberg:2012pjx , the solution of which gave a strong boost to the birth of the inflation theory Starobinsky:1980te ; Guth:1980zm . Because magnetic monopole is deeply connected with the fundamental symmetries of nature at very high energy, the discovery of even a single magnetic monopole would have far-reaching implications [^2] We refer the reader to refs. Bai:2021ewf ; MoEDAL:2021vix ; IceCube:2021eye ; Iguro:2021xsu for some recent interesting attempts at searching for magnetic monopoles.. It is thus intriguing that the recent Amaterasu cosmic ray particle found by the Telescope Array Group TelescopeArray:2023sbd might be interpreted as a magnetic monopole with extremely high energy Cho:2023krz ; Frampton:2024shp .
@@
Although it is possible that cosmic inflation has diluted away magnetic monopoles so that they end with an abundance too small to have observable consequences, it is not necessarily the case Preskill:1984gd ; ParticleDataGroup:2022pth . Relatively light magnetic monopoles might be produced after inflation through various mechanisms Kibble:1976sj ; Zurek:1985qw ; Murayama:2009nj ; Das:2021wei ; Kobayashi:2021des ; Kobayashi:2022qpl ; Maji:2022jzu ; He:2022wwy ; Lazarides:2024niy so that their abundance is not affected by the dilution. Even if for superheavy magnetic monopoles that are indeed affected by inflation, their final abundance still depends on its initial value and the actual amount of dilution and which can be quite model-dependent and not well-constrained Epele:2007ic ; Senoguz:2015lba ; Kephart:2017esj ; Lazarides:2019xai ; Chakrabortty:2020otp ; Vento:2020vsq . Alternative scenarios to inflation could also be envisioned Brandenberger:2009jq . Therefore we choose to be open-minded and inquire what model-independent constraints can be obtained on the cosmological monopole abundance, which is taken as a free parameter determined by some unknown production mechanism in the very early universe. Here by “model-independent constraints” we refer to constraints that can be derived from the electromagnetic and gravitational properties of the magnetic monopole only. Other properties that depend on the model detail (such as catalysis of baryon number violation Rubakov:1982fp ; Callan:1982ah ; Callan:1982au ) will not be assumed.
@@
In this work we shall however demonstrate that the Debye shielding bound suggested by ref. Medvedev:2017jdn cannot be used to obtain meaningful bounds on the cosmological monopole abundance. The reason is that, as we shall see, the concept of Debye shielding simply does not hold for magnetic monopole plasma in the presence of magnetic fields that are not curl-free. A similar issue in the context of solar physics was discussed in Appendix B of ref. Spicer:1984uui , where the authors commented that “Debye shielding will not shield out a transverse electric field.” In our analysis we note that Debye shielding is implicitly based on the requirement of thermal and mechanical equilibrium [^3] We discuss the subtle issue of thermal equilibrium of a magnetic monopole plasma at the beginning of Sec. 2.3. which leads to inconsistencies in the magnetic case with test electric currents.
For non-relativistic magnetic monopoles that do not catalyze baryon number violation, the existing bounds on their cosmological abundance then become quite limited. We revisit the mass density constraint, the bound from direct search in cosmic rays and the conventional Parker bounds. We find that the Andromeda Parker bound which was once proposed to constrain the abundance of magnetic black holes Bai:2020spd , can be employed to constrain the abundance of ordinary magnetic monopoles effectively.
@@ 2.3 Anatomy of the magnetic monopole plasma with a two-fluid model
In order to present a solid argument, let us show that the assumption that Debye shielding is achieved in a magnetic monopole plasma with test electric currents leads to inconsistencies, and thus cannot be used to obtain a meaningful constraint on the cosmological monopole abundance. We start by considering the magnetic monopole plasma using a two-fluid model, with one component fluid composed of magnetic monopoles, and the other component fluid composed of magnetic anti-monopoles. This means that the two fluids can be characterized via the usual fluid equations and equations of state. Such a two-fluid treatment is routine practice in the context of electron-ion plasmas in laboratory and space, while in the case of a magnetic monopole plasma one might worry about the validity of such a fluid approximation. In ref. Medvedev:2017jdn the number of magnetic monopoles within a Debye sphere is estimated to be much larger than unity in parameter space of phenomenological relevance [^6] Using notations and some benchmark values in Sec. 3, we may compute the number of magnetic monopoles within a Debye sphere as $N_{D}\simeq 2.3\times 10^{43}\times\left(\frac{mc^{2}}{10^{16}\,\text{GeV}}\right)^{3/2}\left(\frac{v}{10^{-3}c}\right)^{7/2}\left(\frac{g}{g_{D}}\right)^{-3}\left(\frac{F_{M}}{1.3\times 10^{-16}\,\text{cm}^{-2}\text{s}^{-1}\text{sr}^{-1}}\right)^{-1/2},$ which is certainly much larger than unity in all parameter space of interest., so collective effects associated with the monopoles and anti-monopoles should be significant. On the other hand, the magnetic monopole plasma is found to be highly collisionless Medvedev:2017jdn , since the collision frequency due to magnetic Coulomb scattering is very small. Whether collisionless systems can be treated within a fluid approximation is a subtle issue. Zero collision frequency implies an inifinite mean free path, leading to absurd results of divergent coefficients of viscosity and heat conduction. There are known circumstances in which a fluid approximation is inadequate for capturing the essential physics of such collisionless systems Weinberg:2003ur ; Flauger:2017ged .
In the problem at hand the essence of the fluid approximation is the use of a temperature and an equation of state (based on an isotropic pressure) to characterize the magnetic monopole plasma. Then the fluid equations can be simply obtained from integrating the collisionless Boltzmann equation (and its first velocity moment) over the monopole velocity distribution. The equation of state plays the role of a closure relation so that we do not need to consider higher moments of the collisionless Boltzmann equation Thorne:2017mcp . The notion of a temperature and the associated equation of state seems to hinge on a state of thermal equilibrium that can only be achieved through binary collisions. On the other hand, astrophysical plasmas are often collisionless on scales of interest, and collisionless (astrophysical) plasmas are routinely observed or assumed to exhibit collisional, fluid-like behaviors (e.g. ‘Langmuir’s Paradox’) Bret:2015qia . It is conceivable that a number of mechanisms exist to achieve collisionless thermalization. There can be wave-particle scattering due to plasma instabilities. [^7] A similiar phenomenon is the scattering of cosmic rays by Alfvén waves Thorne:2017mcp . For self-gravitating systems phase mixing and violent relaxation can settle the system quickly into a stationary state 2008gady.book…..B . Though out of thermal equilibrium in the strictest sense, in the context of dark matter physics it is possible Hansen:2004dg ; Ling:2009eh that the dark matter velocity distribution can be described by Tsallis distributions derived from nonextensive statistical mechanics, which permits the notion of a temperature 2013SSRv..175..183L .
As to the magnetic monopole plasma, although the effect of Coulomb collisions between monopoles is negligible, motion of magnetic monopoles is subject to the influence of the environmental gravitational field and wave-particle interaction which are yet to be studied. There are two possibilities regarding the outcome of these collisionless thermalization mechanisms. The first possibility is that they are inadequate to bring the magnetic monopole plasma into a state of quasi-equilibrium [^8] In this work we use the word “quasi-equilibrium” (or sometimes the phrase “thermal equilibrium”) in the sense that the distribution function is approximately Maxwellian, or something like Tsallis distributions that can be derived from nonextensive statistical mechanics, so that the notion of a temperature can be used. In case plasma particles obey Tsallis distributions, Debye shielding in an electron-ion plasma can still be discussed 2014JPlPh..80..341L . that can be characterized by the notion of a temperature and the associated equation of state. Since the concept of Debye shielding hinges on such a state of quasi-equilibrium and the notion of a temperature, in this first possibility Debye shielding cannot be used to yield a meaningful constraint on the cosmological monopole abundance, so the plasma constraint obtained by ref. Medvedev:2017jdn does not apply. The second possibility is that the collisionless thermalization mechanisms can bring the magnetic monopole plasma into a state of quasi-equilibrium that can be characterized by the notion of a temperature and the associated equation of state. In this work, we will however show that even if this second possibility is realized, Debye shielding still cannot be used to yield a meaningful constraint on the cosmological monopole abundance.
Keeping the second possibility in mind, we therefore proceed with the two-fluid model, and write down the Euler equations for the magnetic monopole fluid and the anti-monopole fluid: [^9] Here we neglect terms representing the viscous force. Inclusion of viscous force terms does not alter the conclusion of the argument.
@@
in which v is the fluid velocity associated with the electric currents101010v should be thus related to $\textbf{J}_{e}$ via an equation like Eq. (18), with an additional equation for the continuity of the electric charge. These details are not needed in the following so we do not bother to present them explicitly., and $\sigma$ denotes the electric conductivity which may be a function of the position and time. Finally, there should be equations of state which express the pressure as a function of the number density and temperature:
@@
Similar assumptions are made in the usual treatment of electron-proton plasma in a two-fluid model Thorne:2017mcp . Besides the subtlety discussed at the beginning of this subsection, the first assumption implies that we are effectively working at scales much larger than the mean separation between magnetic monopoles and the magnetic monopole plasma is not very dense so that the monopole annihilation can be neglected.
@@ 3 Revisiting current bounds on the cosmological monopole abundance
    Parker bounds.
All these bounds have been discussed to various extent in previous literature Burdin:2014xma ; Patrizii:2015uea ; Patrizii:2019eud ; Mavromatos:2020gwk ; ParticleDataGroup:2022pth ; Kobayashi:2023ryr . [^11] There exist other interesting bounds that constrain the monopole flux or simply the existence of magnetic monopole in the spectrum of the theory, for example bounds from heavy-ion collision Gould:2017zwi ; MoEDAL:2021vix , neutron star Gould:2017zwi ; Hook:2017vyc , cosmic ray collision with atmosphere Iguro:2021xsu . These bounds however only constrain light magnetic monopoles with $m\lesssim 100\,\text{TeV}$, much lighter than the mass range considered in this work. The emphasis of the discussion here is to examine whether magnetic monopoles may account for all (or the majority) of the dark matter, and the assumptions and uncertainties involved in the analysis. Moreover, motivated by previous studies on magnetic black holes Bai:2020spd ; Ghosh:2020tdu ; Diamond:2021scl , we will improve the existing Parker bound by considering magnetic fields from the Andromeda galaxy.
Bounds on the cosmological monopole abundance can be represented in various manners. In many cases one considers magnetic monopoles (including anti-monopoles) populated with the number density $n_{M}$ and an isotropic average velocity $v$ in some volume defined in a reference frame $S$. Here the reference frame and the volume depend on the type of local measurement under consideration. For example, for direct search in cosmic rays, the reference frame is fixed on the Earth and the volume is associated with the size of the solar system, while for the Parker bound from the galactic magnetic field, the reference frame is fixed on the Milky way galaxy and the volume is associated with the galactic magnetic field region in the Milky way. The local isotropic flux of magnetic monopoles is defined as
$F_{M}=\frac{n_{M}v}{4\pi},$ (43)
Eq. (43) means that if one considers an area element $dS$ with its normal direction n, then during a short time $dt$ the number of monopoles (and anti-monopoles) that cross $dS$ with velocity direction in a small solid angle $d\Omega$ around n is given by $F_{M}dtdSd\Omega$. Therefore the unit of $F_{M}$ is written as $\text{cm}^{-2}\text{s}^{-1}\text{sr}^{-1}$. To gain an intuitive feeling of the flux, we may compute the mean separation $d_{M}=n_{M}^{-1/3}$ of magnetic monopoles as a function of $F_{M}$ and $v$ as
$d_{M}\simeq 288\,\text{km}\times\left(\frac{v}{10^{-3}c}\right)^{1/3}\left(\frac{F_{M}}{10^{-16}\text{cm}^{-2}\text{s}^{-1}\text{sr}^{-1}}\right)^{-1/3},$ (44)
@@
In order to discuss the conversion between $F_{M}$ and $\Omega_{M}$, we introduce $\gamma\equiv(1-v/c)^{-1/2}$ which is the Lorentz boost factor of the magnetic monopoles measured in the reference frame $S$ associated with the local measurement under consideration, and $\rho_{M}=n_{M}mc^{2}\gamma$ which is the local energy density of the magnetic monopoles in the $S$ frame. The conversion between $F_{M}$ and $\Omega_{M}$ is then given by
$F_{M}$ $=\frac{\rho_{\text{crit}}\Delta v}{4\pi\gamma mc^{2}}\Omega_{M},$ (47)
$\Omega_{M}$ $=\frac{4\pi\gamma mc^{2}}{\rho_{\text{crit}}\Delta v}F_{M},$ (48)
@@
The conversion between $F_{M}$ and $\Omega_{M}$ critically depends on the velocity $v$ of the magnetic monopoles. For monopoles clustered with the galaxy ($m\gtrsim 10^{18}\,\text{GeV}/c^{2}$), $v\simeq 10^{-3}c$ which is the virial velocity of the galaxy. Due to the presence of galactic and intracluster magnetic fields lighter magnetic monopoles can be accelerated to higher (and even relativistic) velocities. It is estimated that magnetic monopoles with masses around $10^{16}\textendash\textendash 10^{17}\,\text{GeV}$ can be accelerated to $v\simeq 10^{-2}c$ Medvedev:2017jdn ; ParticleDataGroup:2022pth which is the escape velocity of large galaxy clusters Medvedev:2017jdn , while magnetic monopoles with masses smaller than $10^{11}\textendash\textendash 10^{13}\,\text{GeV}$ are expected to be in the relativistic regime Wick:2000yc ; Burdin:2014xma ; Medvedev:2017jdn ; ParticleDataGroup:2022pth . The dependence of the average monopole velocity on its mass is associated with its complicated acceleration history Wick:2000yc ; Perri:2023ncd and thus cannot be predicted precisely.
To facilitate comparison between different constraints, let us express Eqs. (47) and (48) using various benchmark values (we work in the non-relativistic regime here so that $\gamma\simeq 1$; for $\Delta$ we consider two benchmark values $\Delta=1$ and $\Delta=2.86\times 10^{5}$ so there are two equations each for $F_{M}$ and $\Omega_{M}$)
@@
Here for $B$ we consider only the regular magnetic field Fletcher:2003ec . The Parker bound from magnetic fields of the Andromeda galaxy was employed by ref. Bai:2020spd ; Ghosh:2020tdu ; Diamond:2021scl to constrain the abundance of extremal magnetic black holes. It was then pointed out in ref. Kobayashi:2023ryr that whether such magnetic black holes remain clustered with the Andromeda galaxy is uncertain. Here we use the Andromeda version of the Parker bound to constrain the abundance of ordinary magnetic monopoles with $g=g_{D}$. The boundary between clustered and unclustered monopoles in the case of the Andromeda galaxy is expected to be at $m\simeq 10^{19}\,\text{GeV}/c^{2}$ Kobayashi:2023ryr , again with a transition range spanning $2\sim 3$ orders of magnitude. Therefore somewhat large uncertainty on the overdensity $\Delta$ is involved if one wants to convert the Andromeda Parker bound on the monopole flux at large monopole mass $m\sim 10^{19}\,\text{GeV}/c^{2}$ to a corresponding bound on the monopole density parameter. For much lighter magnetic monopoles with $m\lesssim 10^{19}\,\text{GeV}/c^{2}$ it is likely that they are unclustered with $\Delta\simeq 1$.
@@
$l_{c}\lesssim\frac{v}{2g}\left(\frac{mv}{F_{M}}\right)^{1/2}.$ (65)
Using benchmark values, Eq. (65) can be expressed as [^13] To obtain the numbers, we use $g_{D}\simeq 8.2\times 10^{-7}\,\text{GeV}^{1/2}\,\text{cm}^{1/2}$ in Gaussian units, which can be derived from the Dirac quantization condition $eg_{D}=\hbar c/2$ and $\alpha=\frac{e^{2}}{\hbar c}$.